# PrivaDE: Privacy-preserving Data Evaluation for Blockchain-based Data Marketplaces

### v1

@@ 5.2. The PrivaDE Protocol
PrivaDE securely realizes $\mathcal{F}_{\mathsf{Score}}$ with malicious security.
The proof of Theorem 5.1 will be provided in Appendix C.1.
@@ 6. Blockchain Integration for PrivaDE
After scoring (potentially across multiple contributors), the model owner proceeds with transactions to selected data owners. Transactions can use hash-locked exchanges (Poon and Dryja, 2016):
@@ 7. Evaluation
To assess feasibility, we benchmark the secure protocol with respect to runtime, memory usage, communication cost, and related metrics. Scalability is analyzed by varying both model size and dataset complexity. We further compare our active data selection strategy against entropy-based sampling and the Core-Set method (Sener and Savarese, 2018), evaluating its effectiveness in selecting batches that most improve model performance.
Our open-source implementation is publicly available. We instantiate SMPC using the SPDZ2k protocol within MP-SPDZ (Keller, 2020), employ EZKL (zkonduit, 2023) for ZKP-based inference, and implement the challenge protocol in Circom (Bellés-Muñoz et al., 2023).
We validate on MNIST (LeCun and Cortes, 2010), CIFAR-10 (Krizhevsky, 2009), and CIFAR-100, testing three representative models: LeNet (61,706 parameters), ResNet-20 (272,474 parameters), and VGG-8 (4,506,276 parameters).
@@ 7.1. Practicality of PrivaDE
We implement PrivaDE in a two‐party simulation. Each party executes as a separate process on a single host (32‐core Intel Core i9, 128 GB RAM), communicating over distinct TCP ports. The loss function is always cross‐entropy $\mathcal{L}_{\mathrm{CE}}$; uncertainty is computed as average prediction entropy; diversity is computed as average feature‐wise standard deviation. The representative set is selected via the kk‐center greedy algorithm.
@@
Table 2 reports results of the $\mathsf{CP}$ protocol for representative set verification. Setup time refers to SRS generation, key generation, and program compilation, while online time includes both proof generation and verification. Results show stable online performance across experiments, as $\mathsf{CP}$ runtime is independent of model size; all online times remain below 60 seconds.
The model split procedure (A–B–C) is assumed precomputed and validated (see App. D). Table 3 presents performance of part A inference within malicious 2PC. Online timings are consistently low (sub-20 seconds), with the offline phase dominating due to heavy communication overhead. Since offline time consists primarily of input-independent randomness generation, it can be completed prior to execution.
Inference for parts B and C, executed locally but verified via $\mathsf{CnCZK}$ (Tables 3 and 4), show modest resource requirements: both setup and online costs are low. This demonstrates that cut-and-choose significantly reduces time, memory, and communication overhead while still providing reasonable assurance.
@@
Overall, Table 5 summarizes the end-to-end runtime of the full protocol. While PrivaDE still requires nontrivial resources, it is feasible: smaller models require only about one minute of online time and less than 15 minutes end-to-end. Further optimizations—such as reducing the size of block A or lowering the number of $\mathsf{CP}$ and $\mathsf{CnCZK}$ checks—can trade off security assurance for reduced runtime, depending on application needs.
@@ 7.2. Robustness of the scoring algorithm
The multi‐point scoring algorithm (Algorithm 1) is evaluated against entropy‐based sampling (Settles and Craven, 2008) and core‐set sampling (Sener and Savarese, 2018) on MNIST, CIFAR-10 and CIFAR-100. After normalization and Gaussian blurring, each dataset is split into a pre‐training set (Alice’s initial data), a held‐out test set, and $N$ contributing sets with skewed class distributions to simulate heterogeneity.
Using LeNet for MNIST, ResNet20 for CIFAR-10 and VGG-8 for CIFAR-100 (same setup as above), our framework comprises: representative selection $\mathcal{S}$ via K‐Center‐Greedy (Sener and Savarese, 2018); loss $\ell=\mathcal{L}_{\mathrm{CE}}$; uncertainty $\mathcal{U}$ as Shannon entropy; diversity
@@ 8. Related Works
Song et al. (Song et al., 2021) propose a data evaluation protocol based on uncertainty sampling, later refined by Qian et al. (Qian et al., 2024). Their main advance is an inner-product-based functional encryption scheme that accelerates inference. However, both works adopt the semi-honest adversarial model, which is insufficient for blockchain-based deployments where no trust is assumed between parties.
@@
Tian et al. (Tian et al., 2022) take a different approach by estimating Shapley values via a learned Data Utility Model and then executing custom MPC to perform inference and approximate Shapley scores. This method, however, requires sharing portions of the data owner’s samples and additional labeling to train the utility model; moreover, the end-to-end MPC cost of full-model inference remains high (e.g., $\sim$16 hours for 2,000 CIFAR-10 images in their evaluation), which is ill-suited to marketplaces that demand quick, low-effort data assessment and purchase.
@@
Data pricing. Data pricing is usually concerned with setting a price for data in a standard currency. The research in this area adapts economic principles of pricing and markets to data (Zhang et al., 2023), and is concerned with the data-specific challenges such as duplication and arbitrage (Chen et al., 2019; Ohrimenko et al., 2019). These works take the perspective of the broker and consider the seller and buyer’s market separately. The actual utility of the data to the consumer is assumed to be determined by simple attributes (Koutris et al., 2015).
@@
Game theoretic fair data valuation. This topic emerged with the popularity of machine learning and demand for data (Jia et al., 2019; Ghorbani et al., 2020; Ghorbani and Zou, 2019), and can be seen as an aspect of explainable machine learning (Rozemberczki et al., 2022). The aim of game theoretic data valuation is to determine fair value for data points (or subsets) in a multiple data contributor setting. These works make use of Shapley Value (Rozemberczki et al., 2022) with its Fairness properties to ensure fair valuation for all. Crucially, in these problems the fairness consideration is between different data contributors, as Shapley Value is a mechanism for determining fair share between multiple players. Several associated approaches exist for evaluating data, such as leave-one-out testing (Cook, 1977) and influence function estimation (Koh and Liang, 2017; Datta et al., 2015; Pruthi et al., 2020), with similar objectives. Shapley value and similar techniques tend to be computationally expensive – repeatedly training and testing models on different subsets of data – and require a trusted party to carry out the computation.
@@
Active learning. We borrowed several ideas from this area, such as the use of uncertainty and diversity (Settles, 2009; Dasgupta, 2011). Methods that use both are called hybrid methods. There are algorithms incorporating different scoring values non-linearly (Ash et al., 2020; Yang et al., 2017; Zhdanov, 2019) and linearly (Cardoso et al., 2017). In the standard active learning paradigm, loss is not a decision criterion, because loss computation requires the ground truth labels, which are assumed to be expensive, and the algorithms aim to minimize the use of labels. In (Yoo and Kweon, 2019), a loss predictor is trained simultaneously with the model to inform the active selection process.
@@
Zero knowledge proofs in Machine learning. Researches in this area have been mainly focused on generalizability and efficiency. For generalizability, there are works mapping machine learning operations into suitable zero-knowledge circuit, and translation of non-linear and floating point operations into suitable approximate encryption-friendly circuits (e.g. (Ghodsi et al., 2017)). Other researches in this area have approached the efficiency problem by making use of specific features of models and training algorithms (Angel et al., 2022). However, current zero knowledge proof systems are considered useful mainly for model inference and they are not efficiently supporting the training process. Commonly they affect the model’s accuracy or can be too slow for training even moderate sized models.
@@
Secure MPC Inference Methods. Although zero‐knowledge proofs (ZKPs) protect model weights and guarantee correctness of inference, they do not by themselves ensure privacy of the query input $x$. An alternative is to employ maliciously‐secure multiparty computation (MPC) under a dishonest–majority model, which provides both confidentiality and integrity in a two‐party setting. The most prominent example here is MD‐ML (Yuan et al., 2024), which optimizes cryptographic primitives to minimize online inference latency. While MD‐ML achieves low absolute latency on small models, its performance degrades sharply with complexity––incurring roughly a 40× slowdown from a Linear SVM to LeNet.
Many other maliciously‐secure inference frameworks assume an honest–majority among three or more parties, which is incompatible with two‐party scenarios. Representative systems in this category include Helix (Zhang et al., 2025b), Blaze (Patra and Suresh, 2020), and SWIFT (Koti et al., 2021). Because these require a strict majority of honest participants, their security and performance guarantees do not directly translate to our setting.

### v2

@@ 4.3. The PrivaDE Protocol
PrivaDE securely realizes $\mathcal{F}^{k}_{\mathsf{Score}}$ with malicious security.
The proof of Theorem 4.2 will be provided in Appendix D.1.
5. Practical Optimizations for PrivaDE
In this section, we introduce the design techniques that make PrivaDE’s secure model inference practical in terms of computation and communication.
5.1. Model Distillation
Model distillation for modern neural networks, introduced by Hinton et al. (hinton2015distilling), is widely used to reduce model size and inference cost. The original (teacher) model guides training of a smaller (student) model, typically by matching the teacher’s logits.
Distillation is also common in secure inference (e.g., (splithe)) to lower protocol runtime. Empirical studies, such as (Romero2014FitNetsHF; efficacyKD), showed that student models with over $10\times$ fewer parameters can retain accuracy comparable to the teacher, depending on task and architecture.
In PrivaDE, we include optional local distillation by the model owner prior to data evaluation, where the choice of student model balances accuracy loss and size reduction. Beyond efficiency, distillation can reduce leakage: limited-capacity students cannot faithfully encode all teacher information (efficacyKD). We deliberately select a student size tolerating minor accuracy loss for substantial runtime savings and reduced leakage. Section 7 shows reduced models still support effective data evaluation.
5.2. Split Model
Most ML models, particularly neural networks, compose layers: one layer’s output feeds the next until the final layer. This enables split learning (splitnn), partitioning a network at a cut layer. Typically, the data owner trains the front segment and the model owner trains the back segment, coordinating to reduce data exposure.
In secure inference, this yields split inference (e.g., (splithe; c2pi)): the front segment executes within a secure protocol, while remaining layers run in the clear, reducing cryptographic workload and latency while limiting input leakage to the (carefully chosen) intermediate activations at the cut layer.
In PrivaDE, we also employ split-model inference, but use a three-way split that factorizes $M$ as $C\circ B\circ A$ (Figure 6):
  1. (1)
    Block $A$. Contains the initial layers up to (and including) the first activation. We insert a secret mixer that permutes $A$’s outputs; this mixer is an invertible linear layer or an invertible $1\times 1$ convolution that mixes channels (flow). Block $A$ is evaluated via a maliciously secure 2PC: Alice supplies weights, Bob supplies inputs, and only Bob learns the output of $A$.
  2. (2)
    Block $B$. Comprises the intermediate layers. Its first layer incorporates the inverse mixer to undo $A$’s permutation. The BBCC cut is chosen by the C2PI boundary-search procedure (c2pi). Bob executes $B$ locally after receiving the corresponding weights from Alice.
  3. (3)
The BBCC split follows prior work that balances input-leakage mitigation and runtime. We adopt the C2PI boundary search (c2pi), which probes candidate cut layers by attempting input reconstruction from cut-layer activations; for images, a split is accepted when the structural similarity index with the original input falls below $0.3$, which represents the image not being recognizable by human eye.
The AABB split further reduces 2PC overhead by restricting cryptographic computation to the minimal front segment (up to the first activation), leaving the remainder in $B$. Because $B$’s weights are sent in the clear to Bob, $A$ prevents him from obtaining a self-sufficient front model that yields reusable embeddings on arbitrary inputs.
Finally, we attach zero-knowledge proofs to the evaluations of $B$ and $C$ to enforce correctness and deter cheating; see Sec. 4 for protocol details.
5.2.1. Leakage Analysis.
In our split-model framework, partial leakage of models and activations is unavoidable as it is not protected by zero-knowledge proofs or multiparty computation. Following (c2pi), we therefore adopt an empirically defined client data privacy model to justify and quantify the leakage.
Before executing PrivaDE, both parties must decide whether the resulting leakages are acceptable according to metrics obtained from experimental attacks; if not, they abort the protocol.
Leakage of model $B$ to Bob: Yosinski et al. (featureTrans) show that later layers are more task-specific, while early layers capture generic features. Since neither the weights of $C$ nor the model outputs are revealed, Bob cannot reconstruct the full task model. However, $B\circ A$ may still provide useful embeddings and thus be valuable for fine-tuning on related domains; we therefore need to bound the transferability of $B$.
To repurpose the model, Bob must
  1. (1)
    reconstruct $A$ to map raw inputs into $B$’s input space; and
  2. (2)
    apply transfer learning to $B\circ A$ on new tasks.
We summarize model leakage as a tuple $(\Delta\mathrm{Acc},\textsc{LogME})$, where $\Delta\mathrm{Acc}$ captures the reconstruction quality of $A$ and LogME scores the transferability of $B\circ A$. Alice deems the leakage acceptable or not by comparing $(\Delta\mathrm{Acc},\textsc{LogME})$ to baselines from other marketplace models.
For reconstructing $A$, Bob observes only his inputs and the corresponding outputs of $A$. We assume each contributor issues at most $k$ queries through $A$. A secret mixer (which may vary across runs) applies a random invertible permutation to $A$’s outputs, hindering plug-and-play attacks (e.g., attaching public pretrained fronts to $B$) and collusion by contributors pooling I/O pairs.
Empirical studies (orekondy19knockoff) report that successful extraction typically requires on the order of $10^{4}$ queries; hence security improves when $k\ll 10^{4}$. To evaluate robustness at query budget $k$, we run a reconstruction-under-budget experiment: sample $\{x_{i}\}_{i=1}^{k}$ from $D_{A}$ and train a re-initialized copy $A^{\prime}$ by minimizing
where $\lambda>0$ enforces feature consistency through $B$. After training, we fix $B$ and $C$, replace $A$ by $A^{\prime}$, and measure top-1 test accuracy. Our metric is
A split is deemed robust if $\Delta\mathrm{Acc}$ is substantially negative at practical $k$, indicating that $A$ cannot be effectively reconstructed and $B$ is not easily reused under the stated constraints (cf. the privacy criterion in c2pi).
We also assess the transferability of $B\circ A$ (assuming $A$ is known). Following You et al. (logme), we compute LogME, the log marginal evidence that a Bayesian linear regressor explains labels given features. Concretely, we pass labeled images through $B(A(x))$ to obtain features and evaluate LogME on these embeddings; lower scores imply weaker linear transfer and thus lower leakage risk.
Leakage of activations $B(A(x))$ to Alice: We quantify input leakage via reconstruction quality, training an inverse model that maps $B(A(x))$ back to $x$. For images, we adopt the distillation-based inverse-network attack (DINA) (c2pi) and report the mean structural similarity index (SSIM) on a test set. SSIM, valued in $[0,1]$, measures perceptual similarity in terms of luminance, contrast, and local structure, with higher scores indicating more recognizable reconstructions. We deem a split privacy-preserving when the mean SSIM <0.3<0.3, corresponding to low-fidelity, structurally dissimilar reconstructions.
Justification of the 0.3 SSIM threshold
Following conventions in prior work on model inversion (miaattack; c2pi), reconstructions with SSIM <0.3<0.3 are visually unrecognizable. SSIM is calibrated to human perception, and values around 0.3 fall in the “very poor” quality regime in subjective studies (hanafy2018sarjamming; hsieh2018noreference). We adopt this threshold for direct comparability with C2PI.
5.3. Representative Set Selection
To mitigate this, our protocol requires Bob to extract a much smaller representative subset $D_{R}\subset D_{B}$, and then jointly and securely evaluate Alice’s model only on $D_{R}$. To ensure correctness even if Bob is malicious, we enforce honest sampling via a cut-and-choose subroutine (Sec. 5.3.2). Similar representative-sampling strategies (without security) are standard in batch active learning (sener2018activelearningconvolutionalneural) and related ML tasks (mirzasoleiman2013distributed). While our protocol allows Bob to choose any algorithm that achieves $(d,\delta)$-representativeness, common methods include $k$-center clustering and submodular maximization (nemhauser1978analysis; mirzasoleiman2013distributed).
5.3.1. Dimension Reduction.
Selection involves pairwise distances (Def. 4.1), which becomes expensive as $d$ grows. We apply random linear projections (bingham2001random):
$x^{\prime}=R\,x,\qquad R\in\mathbb{R}^{m\times d},\;m\ll d,$
which preserve distances up to $\varepsilon$-distortion by the Johnson–Lindenstrauss lemma (johnson1984extensions). Alice and Bob jointly sample $R$ via $md$ secure coin flips (blum1981coin), deriving a shared pseudorandom seed. Bob projects locally and runs representative-selection on $\{x^{\prime}\}$, reducing computation and communication while retaining fidelity.
5.3.2. Challenge Protocol for Representative Set Selection
A naive way to securely implement the representative-set algorithm $\mathcal{R}$ is via a maliciously secure two-party computation (2PC). However, because state-of-the-art clustering typically relies on costly Euclidean distance computations, this approach incurs substantial overhead.
As an alternative, we introduce the challenge protocol $\mathsf{CP}$, which does not verify the correct execution of $\mathcal{R}$ directly but instead tests whether the resulting subset satisfies $(d,\delta)$-representativeness. A brief overview of $\mathsf{CP}$ was given in Sec. 4.2, and the full protocol appears in Fig. 10 (App. A.1). Here, we discuss the theoretical guarantees of $\mathsf{CP}$.
Theorem 5.1.
Let $D_{B}$ be a dataset of size $n$, and let $D_{R}\subset D_{B}$ be a purported $(d,\delta)$-representative subset. For any constant $c>0$, if Alice runs the challenge protocol $\mathsf{CP}$ with $|I|=\lceil c\ln n/\delta\rceil$ uniformly random challenges and all challenges succeed, then with probability at least $1-n^{-c}$, $D_{R}$ is $(d,\delta)$-representative.
The proof appears in Appendix B.1. This theorem shows that $\Theta(\ln n/\delta)$ challenges suffice to verify $(d,\delta)$-representativeness with high confidence.
5.4. Audit-Style Cut-and-Choose for ML Inference
While significant research has focused on zero-knowledge proofs (ZKPs) for verifying ML inference (e.g., (zkmlaas; kang2022zkml; zkcnn)), constructing end-to-end proofs remains memory- and computation-intensive.
Taking advantage of the fact that ML models are usually a composition of layers, instead of requiring a full proof, we propose a lightweight protocol that detects incorrect batched inference with tunable detection probability (e.g., 80–95%) by sampling and proving only a small, random subset of local layer transitions. The protocol $\mathsf{CnCZK}$ serves as a drop-in replacement for a NIZKPoK of the language $L_{\textsf{ML-Inf}}$ defined in Sec. 4.3.
$\mathsf{CnCZK}$ allows partial verification of the inference process as follows: The prover first commits to the model parameters, inputs, and all intermediate activations of a forward pass using Merkle trees, and the verifier then samples a random subset of inputs ($D_{S}\subset D_{R}$) and layers $T\subset\{1,\dots,L\}$ to audit. For each sampled position $(i,l)$ (where $i\in D_{S}$, $l\in T$), the prover provides a zero-knowledge proof that the corresponding commitments (of $\theta_{l},a_{l-1,i},a_{l,i}$) and the layer computation $m_{l}(\theta_{l},a_{l-1,i})=a_{l,i}$ is consistent, where $a_{l,i}$ denotes the activation (output) of the model at layer $l$, on the input $i$. The full protocol is specified in Fig. 11 (App. A.2).
In our experiments (Sec. 7), we replace full ZKP verifications with the cut-and-choose protocol CnCZK to explore practical privacy–efficiency trade-offs. When all layers and data points are verified, CnCZK is equivalent to the full PrivaDE protocol in Fig. 5.
5.4.1. Cheating-detection probability
Theorem 5.2.
The proof appears in Appendix B.2. As an illustration, with $N=100$, $L=10$, $\rho=0.10$, $m=25$, and $s=6$, the detection probability exceeds $80\%$. This setting audits only $ms/(NL)=150/1000=15\%$ of all layer evaluations, yielding about an $85\%$ reduction relative to checking every layer of every point.
@@ 6. Blockchain Integration for PrivaDE
After scoring (potentially across multiple contributors), the model owner proceeds with transactions to selected data owners. Transactions can use hash-locked exchanges (poon2016lightning):
@@ 7. Evaluation
To assess feasibility, we benchmark runtime, memory, communication cost, and scalability by varying model size and dataset complexity. We compare our active data selection against entropy-based sampling and Core-Set (sener2018activelearningconvolutionalneural).
Our open-source implementation is publicly available. We instantiate SMPC using the SPDZ2k protocol within MP-SPDZ (mpspdz), employ EZKL (zkonduit_ezkl) for ZKP-based inference, and implement the challenge protocol in Circom (circom).
We validate on MNIST (lecun2010mnist), CIFAR-10 (cifar10), and CIFAR-100, testing three representative models: LeNet (61,706 parameters), ResNet-20 (272,474 parameters), and VGG-8 (4,506,276 parameters).
@@ 7.1. Practicality of PrivaDE
We implement PrivaDE in a two-party simulation. Each party executes as a separate process on a single host (32-core Intel Core i9, 128 GB RAM), communicating over TCP ports. Loss is cross-entropy $\mathcal{L}_{\mathrm{CE}}$; uncertainty is average prediction entropy; diversity is average feature-wise standard deviation. Representative sets use $k$-center greedy.
@@
Parameters reflect typical data valuation exchanges. Dataset size $|D_{B}|$ and model sizes are use-case dependent, while others (representative set size $k$, layers sampled for $\mathsf{CNCZK}$) achieve reasonable security—we require over 85% cheating detection when adversaries corrupt 15% of points.
Table 3 reports $\mathsf{CP}$ protocol results. Setup includes SRS generation, key generation, and compilation; online includes proof generation and verification. Performance is stable across experiments, as $\mathsf{CP}$ runtime is model-size independent; all online times remain below 60 seconds.
The model split (A–B–C) is precomputed and validated (App. E). Table 4 shows part A inference within malicious 2PC. Online timings are consistently low (sub-20 seconds), with offline phase dominating due to communication overhead. Offline consists of input-independent randomness generation, hence completable prior to execution.
Inference for parts B and C, executed locally but verified via $\mathsf{CnCZK}$ (Tables 4, 5), show modest resource requirements. Cut-and-choose verification ($\mathsf{CnCZK}$) significantly reduces time, memory, and communication while providing reasonable assurance.
@@
Table 6 summarizes end-to-end runtime. While PrivaDE requires nontrivial resources, it is feasible: smaller models need only about one minute online and less than 15 minutes total. Further optimizations—reducing block A size or lowering $\mathsf{CP}$ and $\mathsf{CnCZK}$ checks—can trade privacy leakage for reduced runtime.
@@
7.1.1. Scalability Test
We also evaluate how different parameter choices affect the overall online runtime, and hence the scalability, of PrivaDE. Using the MNIST experiment as a representative case, Figure 8 shows how the online runtime scales with $n$ (Bob’s dataset size), $k$ (representative set size), $|I|$ (number of points in the challenge protocol $\mathsf{CP}$), and the ratio of verified computation in $\mathsf{CnCZK}$ for model B (the behaviour for model C is analogous).
Bob’s dataset size $n$ has almost no effect on the online runtime because only the representative set is used in the scoring protocol; increasing the number of data points does not increase the online phase runtime.
All other parameters scale approximately linearly with the online runtime. The representative set size $k$ directly affects the amount of computation required for 2PC inference of model A and the scoring component, while $|I|$ and the verification ratio for $\mathsf{CnCZK}$ increase the runtime of their respective parts of the protocol.
Ultimately, the choice of parameters reflects a security–efficiency trade-off. Verifying more computations on more data points increases both parties’ confidence in the scoring result and reduces the probability that either party can cheat, but the overall online runtime grows proportionally.
Figure 8. Online runtime versus choices of $n$, $k$, $|I|$, and the verification ratio in $\mathsf{CnCZK}$. All parameters except $n$ exhibit approximately linear scaling.
@@ 7.2. Robustness of the scoring algorithm
The scoring algorithm (Algorithm 1) is evaluated against entropy‐based sampling (settles2008analysis) and core‐set sampling (sener2018activelearningconvolutionalneural) on MNIST, CIFAR-10 and CIFAR-100. After normalization and Gaussian blurring, each dataset is split into a pre‐training set (Alice’s initial data), a held‐out test set, and $N$ contributing sets.
To simulate heterogeneous contributors, the contributing sets are randomly skewed: some contain only one labeled class, some consist of a single data point repeated multiple times, and others contain mixed classes with random label distributions. The scoring algorithm is expected to prioritize datasets that most improve the model and thus boost test accuracy.
Using LeNet for MNIST, ResNet20 for CIFAR-10 and VGG-8 for CIFAR-100 (same setup as above), our framework comprises: representative selection $\mathcal{S}$ via K‐Center‐Greedy (sener2018activelearningconvolutionalneural); loss $\ell=\mathcal{L}_{\mathrm{CE}}$; uncertainty $\mathcal{U}$ as Shannon entropy; diversity
@@ 8. Related Works
Song et al. (songDE) propose a data evaluation protocol based on uncertainty sampling, later refined by Qian et al. (qianDE) with an inner-product-based functional encryption scheme. However, both adopt the semi-honest model, insufficient for blockchain deployments.
@@
Tian et al. (privatedv) estimate Shapley values via a learned Data Utility Model and execute custom MPC. However, this requires sharing data samples and additional labeling to train the utility model; moreover, end-to-end MPC cost remains high (e.g., $\sim$16 hours for 2,000 CIFAR-10 images), ill-suited for quick marketplace assessments.
Zhou et al. (zhoupsmi) and Zheng et al. (zheng2023csfl) propose secure data valuation schemes for federated learning setups. Zhou et al. use mutual information to quantify contributions within data coalitions, providing a computationally efficient surrogate for the Shapley value. Zheng et al. build on Federated Shapley Value (FSV) (wang2020principledapproachdatavaluation) and design a homomorphic-encryption-based protocol with two central servers to compute FSV. However, both approaches rely on a semi-honest adversarial model and assume trusted servers.
@@
Data pricing. Data pricing is usually concerned with setting a price for data in a standard currency. The research in this area adapts economic principles of pricing and markets to data (zhang2023survey), and is concerned with the data-specific challenges such as duplication and arbitrage (chen2019towards; ohrimenko2019collaborative). These works take the perspective of the broker and consider the seller and buyer’s market separately. The actual utility of the data to the consumer is assumed to be determined by simple attributes (koutris2015query).
@@
Game theoretic fair data valuation. This topic emerged with the popularity of machine learning and demand for data (jia2019empirical; ghorbani2020distributional; ghorbani2019data), and can be seen as an aspect of explainable machine learning (rozemberczki2022shapleyvaluemachinelearning). The aim of game theoretic data valuation is to determine fair value for data points (or subsets) in a multiple data contributor setting. These works make use of Shapley Value (rozemberczki2022shapleyvaluemachinelearning) with its Fairness properties to ensure fair valuation for all. Crucially, in these problems the fairness consideration is between different data contributors, as Shapley Value is a mechanism for determining fair share between multiple players. Several associated approaches exist for evaluating data, such as leave-one-out testing (cook1977detection) and influence function estimation (koh2017understanding; datta2015influence; pruthi2020estimating), with similar objectives. Shapley value and similar techniques tend to be computationally expensive – repeatedly training and testing models on different subsets of data – and require a trusted party to carry out the computation.
@@
Active learning. We borrowed several ideas from this area, such as the use of uncertainty and diversity (settles2009active; dasgupta2011two). Methods that use both are called hybrid methods. There are algorithms incorporating different scoring values non-linearly (ash2020deepbatchactivelearning; yang2017suggestiveannotationdeepactive; zhdanov2019diverseminibatchactivelearning) and linearly (CARDOSO2017313). In the standard active learning paradigm, loss is not a decision criterion, because loss computation requires the ground truth labels, which are assumed to be expensive, and the algorithms aim to minimize the use of labels. In (yoo2019learning), a loss predictor is trained simultaneously with the model to inform the active selection process.
@@
Zero knowledge proofs in Machine learning. Researches in this area have been mainly focused on generalizability and efficiency. For generalizability, there are works mapping machine learning operations into suitable zero-knowledge circuit, and translation of non-linear and floating point operations into suitable approximate encryption-friendly circuits (e.g. (safetynets)). Other researches in this area have approached the efficiency problem by making use of specific features of models and training algorithms (otti). However, current zero knowledge proof systems are considered useful mainly for model inference and they are not efficiently supporting the training process. Commonly they affect the model’s accuracy or can be too slow for training even moderate sized models.
@@
Secure MPC Inference Methods. Although zero‐knowledge proofs (ZKPs) protect model weights and guarantee correctness of inference, they do not by themselves ensure privacy of the query input $x$. An alternative is to employ maliciously‐secure multiparty computation (MPC) under a dishonest–majority model, which provides both confidentiality and integrity in a two‐party setting. The most prominent example here is MD‐ML (mdml), which optimizes cryptographic primitives to minimize online inference latency. While MD‐ML achieves low absolute latency on small models, its performance degrades sharply with complexity––incurring roughly a 40× slowdown from a Linear SVM to LeNet.
Many other maliciously‐secure inference frameworks assume an honest–majority among three or more parties, which is incompatible with two‐party scenarios. Representative systems in this category include Helix (helix), Blaze (blaze), and SWIFT (swift). Because these require a strict majority of honest participants, their security and performance guarantees do not directly translate to our setting.